### 24 Mar 2026 v1

# A kernel-derived orthogonal basis for spectral functions from Euclidean correlators

Abstract
Spectral functions play a central role in the characterization of a wide range of physical systems, including strongly interacting quantum field theories and many-body systems. Their non-perturbative determination from Euclidean correlation functions constitutes a well-known ill-posed inverse problem and has motivated the development of numerous reconstruction techniques. In this work, we propose a systematic and prior-free framework to represent spectral functions using an orthogonal functional basis derived directly from the kernel of Euclidean two-point correlation functions. By differentiating and partially integrating the Euclidean correlator over restricted time intervals, we identify a set of lattice-accessible constraints and the associated basis functions. These functions can be reorganized into an orthogonal basis with respect to which the spectral function may be approximated in a controlled manner. We demonstrate, using several model spectral functions, that the proposed expansion captures global features of the spectral function and reproduces low-energy transport coefficients with good accuracy when the spectral function is sufficiently smooth. While numerical implementation requires high-precision input for the Euclidean correlator, the present framework is intended as a tool to extract robust constraints and overall structures of spectral functions, rather than as a direct reconstruction method. The approach may thus serve as a complementary ingredient or preprocessing step for existing spectral reconstruction techniques.
keywords:
spectral function, non-perturbative, lattice QCD
1 Introduction
Spectral functions encode essential dynamical information of quantum field theories and many-body systems, including particle excitation spectra, decay rates, and transport coefficients. In lattice QCD and related numerical approaches, such information is accessed indirectly through Euclidean correlation functions. The relation between a Euclidean correlator and the corresponding spectral function is typically expressed as an integral transform with a smooth kernel, whose inversion constitutes a paradigmatic example of an ill-posed problem [1].
A large number of techniques have been developed to address this inverse problem, including Bayesian inference and maximum entropy methods [2, 3, 4, 5, 6], singular value decomposition [7], Padé-based approaches [8], the Backus–Gilbert method [9, 11, 10], and other regularized reconstruction schemes [12, 13, 14, 15, 16, 17]  [^1] For reviews, see, for example, Ref. [1] and references therein.. While these methods have achieved important progress, they often rely on priors, regularization assumptions. Establishing systematic approaches that make minimal assumptions remains a topic of active interest. See for recent progress in this direction, for exmaple, Refs. [18, 19, 20].
In this work, we explore a complementary direction. Instead of aiming at a direct reconstruction of the spectral function, we focus on extracting a set of constraints and global features that are inherent to the Euclidean correlator itself. By differentiating the Euclidean two-point function with respect to Euclidean time and integrating over restricted time intervals, we derive a family of relations that constrain weighted averages of the spectral function. The weight functions emerging from this construction are fully determined by the kernel of the Euclidean correlator and decay exponentially at large frequencies.
We show that these kernel-derived functions can be reorganized into an orthogonal basis, with respect to which the spectral function divided by frequency can be approximated in a systematic expansion. The expansion coefficients are given directly by the aforementioned Euclidean constraints and do not require the introduction of priors. The resulting approximation is designed to capture the overall structure of spectral functions and, in particular, to provide stable access to low-energy quantities such as transport coefficients.
The framework presented here is primarily methodological in nature. Its purpose is not to replace existing spectral reconstruction techniques but rather to provide an additional analytical tool that can be used to test spectral ansätze, generate physically motivated constraints, or supply input to more refined reconstruction methods. To illustrate its properties, we apply the method to several model spectral functions under idealized conditions and show that the approximation could be even very precise for smooth spectral functions. A detailed investigation of its numerical performance with realistic lattice data is left for future work.
Related approaches to the present work include both reconstruction-oriented and structure-oriented methods for spectral analysis. Bayesian inference techniques, such as the maximum entropy method and stochastic analytic continuation, aim at reconstructing spectral functions by incorporating prior information and regularization criteria. In contrast, methods based on kernel decomposition, including the Backus–Gilbert approach and Mellin-transform-based expansions, focus on identifying well-conditioned combinations of Euclidean data that can be reliably constrained.
The framework developed here belongs to the latter category. It does not attempt to reconstruct local features of spectral functions but instead provides a kernel-derived orthogonal basis and a set of lattice-accessible constraints. As such, it is naturally suited to serve as a diagnostic tool or as complementary input to more elaborate reconstruction schemes.
The paper is organized as follows. In sec. 2, we derive a family of relations emerging from the corresponding Euclidean correlation function and identify the pairs of constraint and basis function. In sec. 3, a spectral function is expressed as an expansion in the orthogonal basis constructed from the original basis functions. In sec. 4, our approach is illustrated by introducing a few spectral functions by hand. Finally, possible applications and potential challenges in the practical implementation are discussed in sec. 5.
2 Constraints and basis functions
For simplicity, throughout this paper we focus on the Euclidean two-point correlation function of a Hermitian bosonic operator in an equilibrium system, as typically computed in lattice QCD. Following the notation of [21], the relation between a spectral function $\rho(k^{0},{\bf k})$ and a Euclid two-point correlation function $\Pi(\tau,{\bf k})$ in the $\tau$-$\bf k$} space is given by
$\Pi^{E}(\tau,{\bf k})=$ ${1\over\pi}\int^{\infty}_{0}\!\!\!dk^{0}\,\rho(k^{0},{\bf k})\,\frac{\cosh\left[\left({\beta\over 2}-\tau\right)k^{0}\right]}{\sinh\left({\beta\over 2}k^{0}\right)}\,$, (1)
[…]
Consider differentiating $\Pi^{E}(\tau,{\bf k})$ with regard to $\tau$ $n$ times and then integrating over a certain range of $\tau$. Through this operation on (1) with different values of $n$ and different integration ranges, one can derive a series of equations, which can be seen as the constraints on spectral function. The simplest one would be obtained by simply integrating $\Pi^{E}(\tau,{\bf k})$ from 0 to $\beta$ with $n=0$,
$C({\bf k})=$ $\int_{0}^{\beta}\!\!d\tau\,\Pi^{E}(\tau,{\bf k})=2\,\int_{0}^{\infty}\!\!{dk^{0}\over\pi}{\rho(k^{0},{\bf k})\over k^{0}}\,$, (2)
which is nothing but the zero frequency ($k_{n}=0$) modes and provides a "sum rule" on spectral function [21]. For general $n$ and integration ranges, one obtains
$C_{n}^{(m)}({\bf k})=$ ${1\over Z_{n}^{(m)}}\int_{l_{m}}^{u_{m}}\!\!\!d\tau\,\frac{d^{n}\Pi^{E}(\tau,{\bf k})}{d\tau^{n}}\ =\int_{0}^{\infty}\!\!dk^{0}{\rho(k^{0},{\bf k})\over k^{0}}S_{n}^{(m)}(k^{0})\qquad(n=0,1,2,\cdots)\,$, (3)
where the different integration ranges are distinguishes by the label $m$. Hereafter we focus on the determination of $\rho/k^{0}$ rather than $\rho$. $Z_{n}^{(m)}$ is an arbitrary constant and can be freely chosen. In this work, we choose $Z_{n}^{(m)}$ to normalize the integral of $S_{n}^{(m)}(k^{0})$ over $k^{0}$, i.e.
$\int_{0}^{\infty}\!\!dk^{0}\,S_{n}^{(m)}(k^{0})=1\,.$ (4)
The functions $S_{n}^{(m)}(k^{0})$, which emerges from the kernel of the integrand in (1) through the operation, is given by
$S_{n}^{(m)}(k^{0})=$ ${s_{n}^{(m)}(k^{0})\over Z_{n}^{(m)}}={1\over Z_{n}^{(m)}}{k^{0}\over\pi\sinh\left({\beta\over 2}k^{0}\right)}\int_{l_{m}}^{u_{m}}\!\!\!d\tau\,\frac{d^{n}\,\cosh\left[\left({\beta\over 2}-\tau\right)k^{0}\right]}{d\,\tau^{n}}$
$=$ $\left\{\begin{array}[]{ll}{1\over Z_{n}^{(m)}}{-(k^{0})^{n}\over\pi\sinh\left({\beta\over 2}k^{0}\right)}\left(\sinh\left[\left({\beta\over 2}-u_{m}\right)k^{0}\right]-\sinh\left[\left({\beta\over 2}-l_{m}\right)k^{0}\right]\right)&\quad\mbox{for }n=0,2,4,\cdots\,,\\ {1\over Z_{n}^{(m)}}{(k^{0})^{n}\over\pi\sinh\left({\beta\over 2}k^{0}\right)}\left(\cosh\left[\left({\beta\over 2}-u_{m}\right)k^{0}\right]-\cosh\left[\left({\beta\over 2}-l_{m}\right)k^{0}\right]\right)&\quad\mbox{for }n=1,3,5,\cdots\,.\end{array}\right$. (7)
We call $S_{n}^{(m)}(k^{0})$ the basis function. If one chooses $l_{m}=0$ or $u_{m}=\beta$, the $k^{0}$ integral in (3) is not guaranteed to be finite. By eliminating these points, all basis functions decay exponentially in the large $k^{0}$ region and the integral is made definitely finite.
Note that $C_{n}^{(m)}({\bf k})$ in (3) can be calculated solely from $\Pi^{E}(\tau,{\bf k})$ without any information on $\rho/\omega$. Therefore, one can use these constraints to test spectral functions obtained by some method by calculating both sides of (3) if $\Pi^{E}(\tau,{\bf k})$ is available.
Before going into details, let us make the above quantities dimensionless for later use. Supposing that $\rho$ has mass dimension $d_{\rho}$ and defining the following dimensionless quantities,
$\tilde{k}^{0}={\beta k^{0}\over 2\pi},\ \ \tilde{\omega}={\beta\omega\over 2\pi},\ \ \tilde{{\bf k}}={\beta{\bf k}\over 2\pi},\ \ \tilde{\tau}={\tau\over\beta}\,$, (8)
the relevant quantities in dimensionless unit are written as
[…]
$\tilde{s}_{n}^{(m)}(\tilde{k}^{0})=$ $\beta^{n}\,s_{n}^{(m)}(2\pi k^{0}/\beta)={2\tilde{k}^{0}\over\sinh\left(\pi\tilde{k}^{0}\right)}\int_{\tilde{l}_{m}}^{\tilde{u}_{m}}\!\!\!d\tilde{\tau}\,\frac{d^{n}\,\cosh\left[\left(1-2\tilde{\tau}\right)\pi\tilde{k}^{0}\right]}{d\tilde{\tau}^{n}}\,$, (11)
$\tilde{Z}_{n}^{(m)}=$ $\int_{0}^{\infty}\!\!d\tilde{k}^{0}\tilde{s}_{n}^{(m)}(\tilde{k}^{0})\,$, (12)
$\tilde{S}_{n}^{(m)}(\tilde{k}^{0})=$ $\beta^{-1}S_{n}^{(m)}(2\pi\tilde{k}^{0}/\beta)={\tilde{s}_{n}^{(m)}(\tilde{k}^{0})\over\tilde{Z}_{n}^{(m)}}\,$, (13)
[…]
The possible integration range over $\tilde{\tau}$ is now altered to $0<\tilde{l}_{m}<\tilde{u}_{m}<1$.
There is no guide for efficient choices of $(\tilde{l}_{m},\tilde{u}_{m})$. In this study, they are set to be
$(\tilde{l}_{1},\tilde{u}_{1})=\left({1\over 2},{5\over 8}\right)\,,\ (\tilde{l}_{2},\tilde{u}_{2})=\left({1\over 2},{3\over 4}\right)\,,\ (\tilde{l}_{3},\tilde{u}_{3})=\left({1\over 2},{7\over 8}\right)\,$, (15)
[…]
The basis functions $\tilde{S}_{n}^{(m)}(\tilde{k}^{0})$ thus obtained for $n=0,1,2,3,4$ and $m=1,2,3$ are shown in Fig. 1.
Figure 1: The behaviors of the basis functions, $\tilde{S}_{n}^{(m)}(\tilde{k}^{0})$ for $n=0,1,2,3,4$ and $m=1,2,3$. Different colors represent different $n$.
Although it is not shown in visible way, the peak of each basis function tends to shift to the right and/or to lower as $n$ or $m$ increases.
3 Approximating $\tilde{\rho}/\tilde{\omega}$ in an orthogonal basis
We reorganize the original basis $\tilde{S}_{i}(\tilde{k}^{0})$ ($\tilde{S}_{n}^{(m)}(\tilde{k}^{0})$ with $i=3n+m$) into the orthogonal basis $\tilde{Q}_{i}(\tilde{k}^{0})$,
$\tilde{Q}_{i}(\tilde{k}^{0})=M_{j,i}\,\tilde{S}_{j}(\tilde{k}^{0})\,$, (16)
$\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\tilde{Q}^{*}_{i}(\tilde{k}^{0})\,\tilde{Q}_{j}(\tilde{k}^{0})=\delta_{i,j}\,.$ (17)
Here and in the following, the sum is taken over the repeated indices. Substituting (16) into (17), we obtain
$\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\tilde{Q}_{i}^{*}(\tilde{k}^{0})\,\tilde{Q}_{j}(\tilde{k}^{0})=M_{k,i}^{*}\,X_{k,l}\,M_{l,j}=\delta_{i,j}\,$, (18)
where
$X_{k,l}=\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\tilde{S}^{*}_{k}(\tilde{k}^{0})\,\tilde{S}_{l}(\tilde{k}^{0})\,.$ (19)
Since $\tilde{S}_{k}(\tilde{k}^{0})$ is real, $X$ is a real symmetric matrix and diagonalized by the orthogonal matrix $O$,
$O_{j,i}\,X_{j,k}\,O_{k,l}=\lambda_{i}\delta_{i,l}\,$, (20)
with real eigenvalues $\lambda_{i}$. Then, the real matrix $M$ satisfying (18) is constructed by
$M_{i,j}=O_{i,k}\,(\lambda_{k})^{-1/2}\delta_{k,l}\,O_{j,l}\,.$ (21)
We now assume that $\tilde{\rho}/\tilde{k}^{0}$ is well approximated by the expansion in the orthogonal basis, i.e.
$\left({\tilde{\rho}(\tilde{k}^{0},\tilde{{\bf k}})\over\tilde{k}^{0}}\right)_{\rm approx}=\tilde{c}_{i}(\tilde{{\bf k}})\,\tilde{Q}_{i}(\tilde{k}^{0})\,$, (22)
[…]
$\approx$ $\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\left({\tilde{\rho}(\tilde{k}^{0},\tilde{{\bf k}})\over\tilde{k}^{0}}\right)\,\tilde{Q}_{i}(\tilde{k}^{0})$
$=$ $M_{j,i}\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\left({\tilde{\rho}(\tilde{k}^{0},\tilde{{\bf k}})\over\tilde{k}^{0}}\right)\,\tilde{S}_{j}(\tilde{k}^{0})$
[…]
$\left({\tilde{\rho}(\tilde{k}^{0},\tilde{{\bf k}})\over\tilde{k}^{0}}\right)_{\rm approx}=\$$\tilde{C}_{j}(\tilde{{\bf k}})\,M_{j,i}\,\tilde{Q}_{i}(\tilde{k}^{0})=\tilde{C}_{j}(\tilde{{\bf k}})\,M_{j,i}\,M_{k,i}\,\tilde{S}_{k}(\tilde{k}^{0})$
[…]
Just to demonstrate, the explicit form of the approximation using three basis functions with $n=0$ is given by
$\left({\tilde{\rho}(\tilde{k}^{0},\tilde{{\bf k}})\over\tilde{k}^{0}}\right)_{\rm approx}\ =\$$\ \tilde{C}_{0}^{(1)}(\tilde{{\bf k}})\,\left(565.24269\,\tilde{S}_{0}^{(1)}(\tilde{k}^{0})-864.25344\,\tilde{S}_{0}^{(2)}(\tilde{k}^{0})+324.05787\,\tilde{S}_{0}^{(3)}(\tilde{k}^{0})\right)$
$+\tilde{C}_{0}^{(2)}(\tilde{{\bf k}})\,\left(-864.25344\,\tilde{S}_{0}^{(1)(\tilde{k}^{0})}+1342.2126\,\tilde{S}_{0}^{(2)}(\tilde{k}^{0})-521.80259\,\tilde{S}_{0}^{(3)}(\tilde{k}^{0})\right)$
$+\tilde{C}_{0}^{(3)}(\tilde{{\bf k}})\,\left(324.05787\,\tilde{S}_{0}^{(1)}(\tilde{k}^{0})-521.80259\,\tilde{S}_{0}^{(2)}(\tilde{k}^{0})+221.85975\,\tilde{S}_{0}^{(3)}(\tilde{k}^{0})\right)\,$, (25)
where
$\tilde{S}_{0}^{(1)}(\tilde{k}^{0})\ =\$${2\over\tan[\pi/8]}\,{\sinh[\tilde{k}^{0}\pi/4]\over\sinh[\tilde{k}^{0}\pi]}\,$,
$\tilde{S}_{0}^{(2)}(\tilde{k}^{0})\ =\$${1\over\cosh[\tilde{k}^{0}\pi/2]}\,$,
$\tilde{S}_{0}^{(3)}(\tilde{k}^{0})\ =\$$2\tan[\pi/8]\,{\sinh[3\tilde{k}^{0}\pi/4]\over\sinh[\tilde{k}^{0}\pi]}\,.$ (26)
It would be nice if one could expand the space spanned by the orthogonal basis. We attempted to extend the calculation up to $n=6$, but encountered computational difficulties. As the maximum value of $n$ increases, the size of matrix $X$ defined in (19) grows as $3n\times 3n$. Then, as shown in Figure  2, while the largest eigenvalue of $X$ remains in $O(1)$, the smallest eigenvalue decreases exponentially.
Figure 2: The eigenvalue distribution for the different sizes of matrix $X$ which is $i\times i$.
This phenomena appears to commonly happen to the similar situation [25, 1] and to indicate that including more terms may not work so well since the presence of the vanishing eigenvalue means that one of the basis is not linearly independent of others and including such basis does not extend the orthogonal basis. At the same time, since $1/\lambda_{a}$ appears in (24), very large numbers could appear along the way and cancel out to result in an $O(1)$ number where the diagonalization and constraints need to be increasingly accurate to take care of the rounding errors.
4 Examples
Three spectral functions are introduced by hand to examine how well the approximation works. Since the purpose here is to see the mathematical aspect of the expansion, the ideal case is assumed, i.e. not only the Euclidean correlation function at discrete times but also its interpolation to arbitrary time are assumed to be known with no uncertainty.
Three models for the spectral function examined are
model 1 : ${\tilde{\rho}_{1}(\tilde{\omega},0)\over\tilde{\omega}}=\ {4\over(2+\tilde{\omega}^{2})^{2}}\left[1-{\tanh\left(1-\tilde{\omega}\right)\over 4}{\tilde{\omega}^{2}\over\ln^{2}(2+\tilde{\omega}^{2})}\right]\,$, (27)
model 2 : ${\tilde{\rho_{2}}(\tilde{\omega},0)\over\tilde{\omega}}=\ \left[{0.3\over 1+(\tilde{\omega}/0.71)^{2}}+{0.2\over 1+((\tilde{\omega}-7.1)/2.9)^{2}}+{0.2\over 1+((\tilde{\omega}+7.1)/2.9)^{2}}\right]{2\over 1+(\tilde{\omega}/9.5)^{6}}\,$, (28)
model 3 : ${\tilde{\rho_{3}}(\tilde{\omega},0)\over\tilde{\omega}}=\ {15\left\{1+2\,\tilde{\omega}^{2}(\tilde{\omega}^{2}-{1\over 8})(\tilde{\omega}^{2}-1)(\tilde{\omega}^{2}-{9\over 4})+2\,\tilde{\omega}^{2}\right\}\over 60+\tilde{\omega}^{12}}\,.$ (29)
$\tilde{\rho}_{1}$ is borrowed from [17] with a slight modification to guarantee it is always positive, and inspired from a spectral function related to an "electric field" correlator yielding the momentum-diffusion coefficient of a heavy quark [22, 23, 24]. $\tilde{\rho}_{2}$, inspired in two-dimensional Hubbard model, is taken from Ref. [26], in which the methods based on the maximum entropy, singular value decomposition, sampling, and Padé approximants are compared. $\tilde{\rho}_{3}$ is randomly chosen to have a structure different from $\tilde{\rho}_{1}$ and $\tilde{\rho}_{2}$, and probably it is difficult for many methods to reproduce.
Defining "the average" of $\tilde{k}^{0}$ for each basis function by
$\tilde{\omega}_{n}^{(m)}=$ $\int_{0}^{\infty}\!\!d\tilde{k}^{0}\,\tilde{k}^{0}\tilde{S}_{n}^{(m)}(\tilde{k}^{0})$ (30)
and calculating the constraints $\tilde{C}_{n}^{(m)}(0)$ for each of model, we obtain a set of points $(\tilde{\omega}_{n}^{(m)},\,\tilde{C}_{n}^{(m)}(0))$. Let us emphasize that according to our assumption the constraints $\tilde{C}_{n}^{(m)}$ can be calculated solely from the accurately determined Euclidean correlation function without knowing $\tilde{\rho}$.
For example, the first three constraints in each model are calculated to be
$\mbox{model 1 :}\ \tilde{C}_{0}^{(1)}(0)=0.72052116\,,\ \tilde{C}_{0}^{(2)}(0)=0.65470817\,,\ \tilde{C}_{0}^{(3)}(0)=0.48957291\,$, (31)
$\mbox{model 2 :}\ \tilde{C}_{0}^{(1)}(0)=0.51257899\,,\ \tilde{C}_{0}^{(2)}(0)=0.47509899\,,\ \tilde{C}_{0}^{(3)}(0)=0.39432121\,$, (32)
$\mbox{model 3 :}\ \tilde{C}_{0}^{(1)}(0)=0.41800620\,,\ \tilde{C}_{0}^{(2)}(0)=0.43460071\,,\ \tilde{C}_{0}^{(3)}(0)=0.42847991\,$, (33)
Three sets of data points are plotted in Fig. 3 together with the spectral functions of each.
Refer to caption Refer to caption Refer to caption
Figure 3: The models for the spectral function (solid curve) and the constraints obtained from it (dots) for the models $\tilde{\rho}_{1}$ (left), $\tilde{\rho}_{2}$ (middle) and $\tilde{\rho}_{3}$ (right).
They roughly follow the input spectral function in either cases although one can not guess true input only from them.
Substituting the numerical values of the constraints and the basis functions into (24), we obtain the approximated spectral functions shown in Figs. 4, where we have defined the relative difference by
$\tilde{\Delta}={{\tilde{\rho}_{i}\over\tilde{\omega}}-\left({\tilde{\rho}_{i}\over\tilde{\omega}}\right)_{\rm approx}\over{{1\over 2}\left[{\tilde{\rho}_{i}\over\tilde{\omega}}+\left({\tilde{\rho}_{i}\over\tilde{\omega}}\right)_{\rm approx}\right]}}\,.$ (34)
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: The approximated $\tilde{\rho}/\tilde{\omega}$ and the input (black curve) for the model 1 (top), 2 (middle) and 3 (bottom).
In the case of model 1 (see the top panels of Fig. 4), the approximation turns out to successfully reproduce the input in terms of only three basis functions with $n=0$. It means that the approximation
$\left({\tilde{\rho}_{1}(\tilde{k}^{0},0)\over\tilde{k}^{0}}\right)_{\rm approx}=0.08548\,\tilde{S}_{0}^{(1)}(\tilde{k}^{0})+0.5843\,\tilde{S}_{0}^{(2)}(\tilde{k}^{0})+0.4787\,\tilde{S}_{0}^{(3)}(\tilde{k}^{0})\,$, (35)
which is obtained by substituting (31) into (25), is already very accurate. But it should be noted that each coefficient in (35) is the result of the cancellation among large values of $O(10^{2})$ or $O(10^{3})$ and hence the accuracy in the practical calculation will depend on the errors of $\tilde{C}_{n}^{(m)}$ and the correlation among them. $\Gamma=\lim_{\tilde{\omega}\to 0}{\tilde{\rho}(\tilde{\omega})/\tilde{\omega}}$ is often related to a transport coefficient and is also well reproduced better than 2 % for the expansion using only three basis functions.
For the case of the model 2 (see the middle panels of Fig. 4), it turns out that including the nine basis functions corresponding to $n\leq 2$ yields very successful approximation and $\Gamma$ is reproduced within 12 %, 1 %, 0.3 % for $n=0$, $n\leq 1$ and $n\leq 2$, respectively. Unfortunately, again the high-precision numerical results for $\tilde{C}_{n}^{(m)}$ are required in a naive application to practical use. It is interesting to compare our result with those based on four different methods adopted in Ref. [26]. Although the prerequisites are different and hence the totally fair comparison is not possible, our method with $n\leq 2$ provides the best approximation.
For the model 3 (see the bottom panels in Fig. 4), even after including 15 basis functions, i.e. $n\leq 4$, the approximation is not as good as those for the previous two cases, and in particular it fails to reproduce the rapid fluctuation around $\tilde{\omega}=1.5$. Since the basis functions only contain terms like $(\tilde{k}^{0})^{n}$ or $e^{-\tilde{k}^{0}\,\pi}$, the approximation will fail to reproduce the behavior more rapid than those. However, the approximation can be still useful for $\Gamma$. The constraints with $n=0$ are sensitive to the behavior of the spectral function around $\tilde{\omega}=0$ as seen from Fig. 1. Thus, $\Gamma$ can be reproduced relatively well in any cases. Indeed, it is reproduced within 8 % using only $n=0$ constraints for the model 3.
5 Summary
In this work, we have introduced a systematic framework to analyze spectral functions using an orthogonal functional basis derived directly from the kernel of Euclidean two-point correlation functions. By differentiating and partially integrating the Euclidean correlator over restricted time intervals, we constructed a set of lattice-accessible constraints and associated basis functions that decay exponentially at large frequencies. These functions can be reorganized into an orthogonal basis with respect to which the spectral function divided by frequency can be approximated in a controlled expansion.
We have illustrated the properties of this expansion using several model spectral functions under idealized conditions. The results demonstrate that global features of spectral functions, as well as low-energy quantities such as transport coefficients, can be reproduced with good accuracy using only a limited number of basis elements, provided the spectral function does not exhibit rapid oscillatory behavior. At the same time, the analysis reveals inherent numerical challenges arising from the exponentially decreasing eigenvalues, which amplify the sensitivity to statistical and systematic uncertainties in practical applications.
For these reasons, the present framework is not intended as a stand-alone method for the direct reconstruction of detailed spectral functions from lattice data. Instead, it should be viewed as a diagnostic and preprocessing tool that extracts robust, kernel-determined information from Euclidean correlators. Potential applications include testing spectral ansätze, providing physically motivated constraints, or supplying controlled input to more elaborate reconstruction approaches such as Bayesian or regularized inversion methods.
Future work will focus on assessing the numerical performance of this framework with realistic lattice data, including the impact of finite statistics, finite lattice spacing, and limited Euclidean-time resolution. Further developments may also explore optimized choices of integration ranges and strategies to improve numerical stability. We expect that, in these roles, the present approach can contribute a useful methodological element to the broader toolbox for spectral analyses in lattice QCD and related fields.
Acknowledgments
This work is supported in part by JSPS KAKENHI Grant-in-Aid for Scientific Research (No.22K03645).